OpenStudy (anonymous):

A biologist is comparing the growth of a population of flies per week. She has devised an equation to solve for which day (x) the bullfrg would be able to eat the entire population. The equation is 2^x=3x-1. However, she has observed that the bullfrog can not eat more than seven flies in one week. Explain to the biologist how she can solve this on a graph using a system of equations. Identify any possible constraints to the situation.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

she can eat no more than 7 flies per week, so \(\normalsize\color{blue}{ y≤7x\LARGE\color{white}{ \rm │ }}\), where is the number of weeks, and y is the number of flies. \(\normalsize\color{blue}{ \LARGE\color{white}{ \rm │ }}\) Are you sure that the equation is, \(\normalsize\color{blue}{ 2^x=3x-1\LARGE\color{white}{ \rm │ }}\) ?

OpenStudy (anonymous):

Yes, I am

OpenStudy (solomonzelman):

But the only possible solution to that (the \(\normalsize\color{blue}{ 2^x=3x-1\LARGE\color{white}{ \rm │ }}\)) equation is 1.

OpenStudy (solomonzelman):

I am not sure how to go about \(\normalsize\color{black}{ 2^x=3x-1\LARGE\color{white}{ \rm │ }}\)

OpenStudy (anonymous):

I'm positive that's it... are you using the other equation we came up with as a system?

OpenStudy (solomonzelman):

So perhaps as phi said we just need it for one 1.

OpenStudy (solomonzelman):

one week*

OpenStudy (solomonzelman):

I am not sure about this, maybe @phi @IMStuck since they are already here.

OpenStudy (anonymous):

If we solve the given equation, and you got 1, then perhaps we're supposed to substitute the 1 in for x in the equation we created, which would allow us to solve for y. Right?

OpenStudy (phi):

This question has been asked before. I'll look for the post.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Thank you!

OpenStudy (phi):

http://openstudy.com/study#/updates/5372fefbe4b0c1175e755dc4 ganesh posted a graph showing the two possible solutions.

OpenStudy (phi):

The constraint bullfrog can not eat more than seven flies in one week. probably means the frog can eat 7 flies on the first day (and none after that for a week), or 1 per day, or any combination , as long as it's less than 7 for that week. the plot shows x in days. to show the constraint, plot the line y= 7. No solution (in the first week, i.e. x<8) is allowed to be above 7

OpenStudy (anonymous):

I think I understand...based on the problem, I'm supposed to finding the values for x, correct?

OpenStudy (phi):

we solve the problem by finding where the two curves intersect, and then check that these points meet the constraint. We find two possible solutions (3,8) eat 8 flies on the 3rd day or (1,2) eat 2 flies on the first day only the second meets the constraint of 7 or fewer flies can be eaten in the first week.

OpenStudy (anonymous):

That makes so much more sense. Thank you for explaining it to me!

OpenStudy (phi):

yes. She has devised an equation to solve for which day (x) so you find x. you also find the corresponding y (to make sure we meet the constraint i.e. y value is less than 7 )

OpenStudy (anonymous):

Thank you so much for the help!!

OpenStudy (phi):

yw